You invest $8000 in two accounts paying 8% and 5% respectively. If the interest at year's end is 490, how much of that amount did you invest at each rate?
Answer by mananth(16949)

$8000 in two accounts paying 8% and 5%
Total interest = $490
..
let $ x be invested at 8%
8000-x will be invested at 5%
..
0.08x+0.05(8000-x)=490
0.08x+400-0.05x=490
0.03x=90
x=3000
$ 3000 at 8%
$5000 at 5 %
CHECK
0.08*3000+0.05*5000 = 490